OpenStudy (anonymous):

Venus has a mass of 4.86 1024 kg and a radius of 6.05 106 m. (a) What is the acceleration due to gravity on Venus? -------m/s2 (b) How much would a 5.22-kg person weigh on this planet? ------N

OpenStudy (anonymous):

i think you need the gravitational constant for venus. a = G(4.861024)/ (6.05106)^2 a = GM/ r^2 for part b, f = ma, multiply the acceleration from part "a" by 5.22 to get that person's weight in newtons